A body of mass 50 kg is projected vertically upwards with velocity of 100 m/s. 5 seconds after this body breaks into 20 kg and 30 kg.If 20 kg piece travels upwards with 150 m/s then the velocity of other block will be-?

Grade:11

1 Answers

Khandavalli Satya Srikanth
10 years ago

conservatoin of momentum states that "For a collision occurring between object 1 and object 2 in an isolated system the total momentum of the two objects before the collision is equal to the total momentum of the two objects after the collision. That is, the momentum lost by object 1 is equal to the momentum gained by object 2".

velocity of body before breaks is

V=100-9.8*5=51m/sec

Then from conservation of momentum

51*50=20*150+30*VB

VB=-15m/sec

negative sign indicates second body travels downwards with velocity 15m/sec